OpenStudy (anonymous):

if a ball is dropped from 160 cm and rebounds to 120 cm on the first bounce how high will it be one the 2nd and 5th and nth bounce

OpenStudy (aum):

It will be a geometric progression. The common ratio will be 120/160 = 3/4 160, 120, 120 * 3 /4, ....

OpenStudy (anonymous):

2nd bounce=90, 5th=37.97, nth=160(0.75)n